OpenStudy (anonymous):

identify the open intervals where the function f(x)=-5x^2+x+4 is increasing and/or decreasing

OpenStudy (anonymous):

Do you have calculus knowledge? Differentiation in particular.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Cool :) So we will use that. 1st step get the derivative of the function. Go :)

OpenStudy (anonymous):

1-10x

OpenStudy (anonymous):

Cool, now make it equal to 0 and solve the equation

OpenStudy (anonymous):

ok i got 1/10

OpenStudy (anonymous):

Nice, do you know what happens at x=1/10 with this function?

OpenStudy (anonymous):

Where the derivative is =0

OpenStudy (anonymous):

it decreases?

OpenStudy (anonymous):

Nope, the derivative of a function is nearly the same as the rate of change. So when the derivative is 0, the rate of change of the function is 0. It is neither decreasing nor increasing.

OpenStudy (anonymous):

ok so would the answer be infinity?

OpenStudy (anonymous):

No. But we have found the point where it changes behavior. f(x)=-5x^2+x+4 This is an upside down parabola. So in this interval it is increasing \[(-\infty,1/10)\] and \[(1/10 , \infty)\] it is decreasing.

OpenStudy (anonymous):

ok thank you so much!